# Challenges and recommendations for Electronic Health Records data extraction and preparation for dynamic prediction modelling in hospitalized patients - a practical guide

### v1

@@ Abstract
Dynamic predictive modeling using electronic health record (EHR) data has gained significant attention in recent years. The reliability and trustworthiness of such models depend heavily on the quality of the underlying data, which is largely determined by the stages preceding the model development: data extraction from EHR systems and data preparation. We list over forty challenges encountered during these stages and provide actionable recommendations for addressing them. These challenges are organized into four categories: cohort definition, outcome definition, feature engineering, and data cleaning. This list is designed to serve as a practical guide for data extraction engineers and researchers, supporting better practices and improving the quality and real-world applicability of dynamic prediction models in clinical settings.
@@ 1 Background
Predictive modelling using electronic health record (EHR) data has grown over recent years. EHRs provide detailed longitudinal patient data, enabling dynamic (or continuous) diagnosis or prognosis prediction of various outcomes during a patient’s hospitalisation. These prediction models hold the potential to enhance patient outcomes by detecting risks in real-time, enabling timely interventions, and supporting clinical decision-making. Recent breakthroughs in machine learning and deep neural networks (Vaswani, 2017; Lee et al., 2019) have demonstrated their applicability (Tomašev et al., 2021, 2019; Lee et al., ). The growth in EHR-based predictive modelling is accompanied by public access to large-scale health datasets, the development of standards for data formats, tools for data preparation (or “data wrangling”), and data quality assessment frameworks.
The prediction model development is facilitated by the preceding stages of data extraction and data preparation (also sometimes referred to as data preprocessing). Data extraction entails retrieving and organizing the raw EHR data from complex relational databases into a structured format and uses an ETL process for extracting, transforming and loading the data into a data warehouse. Standards such as the Observational Medical Outcomes Partnership Common Data Model (OMOP CDM) (Hripcsak et al., 2015; Park et al., 2023) facilitate structure and terminology consistency for data extracted for research purposes. Publicly available (upon request) EHR extractions, mostly focused on intensive care unit (ICU) or emergency admissions (Johnson et al., 2016, 2023; Pollard et al., 2018; Thoral et al., 2021), facilitate open and reproducible research for predictive models. Many hospitals also extract their EHR data to conduct research and to build prediction models with practical applicability (Chishtie et al., 2023). This diverse range of EHR datasets, whether ICU-specific or hospital-wide, publicly accessible or proprietary, serves as the foundation for a growing body of predictive modelling research.
Data preparation focuses on transforming the extracted data into a usable format for predictive modelling, often a structured “predictor matrix”. Following extraction, data are processed as a first step in “pipelines” that facilitate the model building. Toolsets for data preparation like tidyverse (Wickham et al., 2019) and tidymodels (Kuhn and Wickham, 2020) in R, or pandas (McKinney et al., 2011) and skrub (Lemaitre et al., 2024) in Python aid researchers in the complex task of transforming the extracted EHR data into model-ready formats. Open code frameworks generalizable to a variety of data formats, such as Clairvoyance (Jarrett et al., 2023) and FIDDLE (Tang et al., 2020), or specific to a public extraction, such as the MIMIC extractions (Gupta et al., 2022; Wang et al., 2020), aim to establish reproducible workflows for EHR data preparation.
As extraction, preparation and modelling toolsets and frameworks for EHR data evolve, so too does the demand for high-quality data. A quality assessment framework for EHR data has been proposed by Weiskopf and Weng (Weiskopf and Weng, 2013) for a broad spectrum of clinical research purposes beyond prediction models. METRIC (Schwabe et al., 2024) is a newly proposed framework tailored for prediction studies. Toolsets for OMOP CMD EHR data exploration and quality assessment have been developed as part of the Observational Health Data Sciences and Informatics initiative, such as Achilles (DeFalco et al., 2023) and Data Quality Dashboard (Blacketer et al., 2021), which facilitate more than 3500 data quality checks. Data quality assessments (DQAs) are performed by researchers using EHR data either in a structured manner, referring to DQA dimensions (Lewis et al., 2023; Miao et al., 2023; Thuraisingam et al., 2021; Nobles et al., 2015) or, often, in an unstructured manner in studies in which researchers describe the challenges, successes or “lessons learned” faced during a project (Ferrao et al., 2016; Sauer et al., 2022; Arbet et al., 2021). Generally, data quality issues (or artifacts) are seen as a given to the researcher to deal with, and recommendations for mitigating the quality artifacts are missing. The fact that the extraction process can also introduce artifacts is less often recognized (Honeyford et al., 2022). Data preparation concerns and recommendations are more often described in data pipelines protocols for time-sensitive data preparation (Tomašev et al., 2021). Few studies emphasize the collaboration between researchers and extraction engineers and the iterative process of the data extraction and cleaning (Honeyford et al., 2022; Lewis et al., 2023; Miao et al., 2023), while both steps (extraction and preparation) are crucial to ensuring the final dataset is representative for the prediction task.
The literature addressing data quality challenges during EHR data extraction and preparation for predictive modelling is scattered across various article types: data quality frameworks, data quality assessment studies, “lessons learned” articles, data extraction standards, data quality toolsets and data preparation pipeline frameworks. Often, the focus is on data cleaning, rather than the entire workflow: cohort definition, outcome definition, feature engineering, and data cleaning. Protocols that cover all stages for dynamic prediction models development emphasize avoiding temporal leaks and defining the outcome dynamically throughout a patient’s stay (Tomašev et al., 2021). Despite these varied contributions, a comprehensive overview covering challenges and recommendations for data extraction and preparation for prediction models with clinical applicability is lacking.
@@ 2 Objective
This article provides a comprehensive list of challenges encountered during data extraction and preparation using electronic health record (EHR) data for developing dynamic prediction models for clinical use. It further proposes recommendations with actionable insights, intended for improving the quality of research and the practical applicability of clinical predictive models. This list is intended as a hands-on resource for data extraction engineers and researchers to consult during the extraction and preparation process. We focus on single-hospital structured data, extracted from the ICU system and hospital-wide EHR software. We focus on medium to large-scale extractions and we cover standardized (e.g.: OMOP CMD) and non-standard, proprietary extractions.
@@
The section “EHR Data Flow from Data Collection to Model Building” explains the typical trajectory of data, for both model building and clinical implementation, setting the stage and providing context for the detailed list of challenges that follows. The section “Challenges and Recommendations” lists the challenges and actionable recommendations, which represent the core of the article. These are categorized into four groups: cohort definition, outcome definition, feature engineering, and data cleaning. The “Discussion and Conclusion” section summarizes the recommendations and reflects on their broader implications.
@@ 3 EHR data flow from data collection to model building
The data flow from the patient’s bed to the prediction model building follows three stages: data collection (a), data extraction (b) and data preparation (c) (Figure 1, A). Patient data are either manually entered in EHR software modules or collected by devices and stored in one or multiple databases (a). EHR data are then extracted from relational databases in a simplified (denormalized) format and typically stored in a data warehouse as multiple “base tables” per category, each capturing different aspects of patient data and healthcare events. Examples of base tables naming as per OMOP CDM are: PERSON, DRUG_EXPOSURE, DEVICE_EXPOSURE, CONDITION_OCCURRENCE, MEASUREMENT, NOTE, OBSERVATION, but naming and granularity of extraction can vary for non-standard extractions (b). Researchers building a prediction model either have access to the data warehouse or get a copy of all tables or, under specific ethical considerations, a subset of the base tables or a subset of the patients in the data warehouse (e.g.: admissions during a specified period of interest for patients undergoing mechanical ventilation). Further, they process the extracted data and bring it in a format on which a prediction model can be built (c).
@@
At model implementation time in clinical practice in the EHR software (Figure 1, B), a trigger (e.g.: update of lab results) or a scheduled task (e.g.: every 24 hours) will initiate the request for a prediction. Data are already collected (a) for the patient in the EHR database(s). The same logic as for extraction (b) is reproduced (typically by reusing the queries or code used for extraction). Using the data packed in a specific format (e.g.: json, xml), the prediction service (typically using a REST API interface for communication) is invoked. Data exchange between the EHR and the prediction service is generally performed using the FHIR (Fast Healthcare Interoperability Resources) standard. The prediction service will further prepare the data (c), invoke the model and return a prediction (and additional information, if foreseen) to the EHR software which will present it to users in form of alerts or patient flags within the patient’s chart. The process is typically logged in the EHR system for monitoring purposes. In summary, at real-time prediction, the data collection (a) happens implicitly and is part of the normal clinical flow; data extraction (b) and data preparation (c) are identically reproduced as for model building.
@@
Figure 1: Data flow for model building pipeline (A) and model implementation (B). Two databases are exemplified as data sources, EHR and ICU, although multiple other sources might be used in the hospital’s flow and for data extraction. EHR = Electronic Health Record; DB = database; ICU = Intensive Care Unit; REST API = RESTful application programming interface.
@@
We address some specific aspects for each of the processes of collection, extraction and preparation, which represent transition phases from one data format to another:
@@
    Data collection The EHR database will not reflect with maximum accuracy the “true state” of the patient. First, it will suffer of incompleteness, as not all possible markers and observations can be collected for all patients at all times. The decisions with regards to what data are collected (e.g.: which laboratory tests are ordered and performed) is highly dependent on the patient conditions and on the hospital procedures. Sufficient data are though collected to support the patient’s clinical follow-up and treatment. From this perspective, EHR data differ vastly from data collected for clinical trials, where researchers specify the measurements, measurement methods and collection procedure. Second, nurses and clinicians might have slightly more information than the data collected in the system, either from patient conversations or organizational knowledge. Considering that we do not cover extraction of text notes and reports, tabular data only will always suffer of a level of incompleteness. Nevertheless, tabular data might prove sufficient for specific prediction tasks. Third, both manual data entering and data collected by devices can be at times error-prone, software can have bugs and data recording procedures in the system which will affect the granularity of observed data will change over time. These considerations, which can be summarized as data completeness, correctness and currency, have to be carefully thought by researchers to assess if EHR data are fit for the prediction goal.
@@
    Data extraction The extracted data might not correctly reflect the EHR database. Although data extraction generally aims for completeness of all clinically relevant data, the extraction process can introduce undesired artifacts or can have its own limitations. Whenever extraction logic bugs are detected, these should be corrected and safeguarded through unit and/or integration tests. For one-time or custom data extractions, close collaboration between researchers and data engineers will enhance the quality of the data extraction. Mature extraction platforms will generally be less error prone and researchers can consider the maturity of the extraction platform as a factor influencing the time they will spend on data preparation. The extraction process is typically carried out by a data extraction engineer, data integration developer, data warehouse engineer or ETL developer representing the EHR vendor or the hospital IT department. Hereafter, we refer to this role as the data extraction engineer. They work closely with EHR software developers, hospital IT staff, and sometimes clinical personnel to understand database structures, document extracted data, and perform clinical concept and terminology mapping.
@@
    Data preparation The prepared data might not correctly reflect the extracted data. Undesired artifacts can be introduced by feature engineering or data cleaning. Good documentation of the extraction format, as well as incorporation of clinical knowledge will ensure that the data and the features are not misinterpreted. Coding errors can always occur. Time-sensitive data poses an additional challenge to ensure no temporal leaks are introduced by mistake. Outcome leaks and test-train leaks can also occur if data are not preprocessed carefully. A good practice is to separate train and test sets first, and apply the data preprocessing separately. A modular organization of the code will facilitate unit testing and easy modifications without introducing new errors. The more complex the data preparation, the more error prone it becomes. We recommend that the whole code is unit tested. We generally recommend against manual corrections that cannot be programmatically reproduced. It is estimated that this phase (which generally includes collaboration with data extraction engineers and clinical experts), takes up to three months for an entire team (Tomašev et al., 2021), but such estimations will be highly dependent on the maturity of the extraction process and of the team’s experience with EHR data. The data preparation is performed by a researcher, data scientist, statistician or machine learning engineer, which we will refer to as researcher for the remainder of the paper.
Each stage of processing the data has its own challenges and can introduce new problems, widening the gap between the patient’s state and the data used by the prediction model and ultimately impacting the model building (suboptimal model), model evaluation (biased performance metrics) or model implementation in clinical practice.
@@ 4 Challenges and Recommendations
We list common problems that can be avoided or alleviated and recommendations for mitigation strategies for problems originating in the data collection process (a), the data extraction process (b) and the data preparation process (c). We also focus on problems that can impede the identical reproduction of the extraction and preparation at clinical implementation time. We assume that both the extraction and data preparation processes are reusable or identically reproducible for implementation in clinical practice. We have categorized the challenges and recommendations into four groups: (1) cohort definition (and inclusion/exclusion criteria), (2) outcome definition, (3) feature engineering, and (4) data cleaning, and each group contains problems originating in the collection, extraction or preparation process.
Our insights are drawn from a selective literature review, as well as our experience with various EHR data extractions. We provide mapping of the items to both the Weiskopf and Weng framework (Weiskopf and Weng, 2013) and the METRIC framework (Schwabe et al., 2024), whenever applicable. Weiskopf and Weng do not include dimensions covering the cohort representativeness and completeness of the extracted features, which are important in prediction settings. We will use the Completeness dimension to refer to completeness of data values, as in the original definition, as well as completeness of the cohort and of the extracted features (our extension).
While we aimed to make these challenges as generic as possible, we recognize that specific issues are often unique to individual projects and may not apply to all prediction tasks. We advise users to assess the impact of each listed challenge in their specific context. Similarly, the recommendations might not always be universally applicable and depend on the project context. Our guidance remains pragmatic; at times, the best approach may be to “leave-as-is” to avoid the risk of overcorrection, which can backfire. Following the fitness-for-use principle (Miao et al., 2023), we encourage readers to remain pragmatic and address only the issues relevant to their cohort.

### v2

@@ Abstract
Dynamic predictive modelling using electronic health record (EHR) data has gained significant attention in recent years. The reliability and trustworthiness of such models depend heavily on the quality of the underlying data, which is, in part, determined by the stages preceding the model development: data extraction from EHR systems and data preparation. In this article, we identified over forty challenges encountered during these stages and provide actionable recommendations for addressing them. These challenges are organized into four categories: cohort definition, outcome definition, feature engineering, and data cleaning. This comprehensive list serves as a practical guide for data extraction engineers and researchers, promoting best practices and improving the quality and real-world applicability of dynamic prediction models in clinical settings.
@@ 1 Background
Predictive modeling using electronic health record (EHR) data has become increasingly important in enhancing patient outcomes through real-time risk detection and timely interventions. However, the effectiveness of these models is heavily reliant on the quality and structure of the underlying data, which are influenced by the processes of data extraction and preparation. While recent advancements in artificial intelligence and machine learning (Vaswani, 2017; Lee et al., 2019) have shown promise in this area (Tomašev et al., 2021, 2019; Lee et al., 2020), significant challenges remain in ensuring that the data used for model development are representative and reliable.
Data collection is part of the routine hospital workflow. Data extraction entails retrieving and structuring raw EHR data from hospital databases using an ETL (extract, transform, load) process, with standards like Observational Medical Outcomes Partnership Common Data Model (OMOP CDM) facilitating structure and terminology consistency. Publicly available ICU and emergency admission datasets (Pollard et al., 2018; Thoral et al., 2021; de Kok et al., 2023; of Us Research Program Investigators, 2019), such as MIMIC (Johnson et al., 2016, 2023), support reproducible research, while hospitals also extract private, not publicly shared data for predictive modeling. Data preparation transforms extracted data into a structured format for predictive modeling, often using R or Python pipelines. Open-source frameworks, whether generalizable (Jarrett et al., 2023; Tang et al., 2020) or specific to MIMIC (Gupta et al., 2022; Wang et al., 2020), aim to establish reproducible workflows for EHR data preparation. Data quality issues, such as inconsistencies and artifacts introduced during extraction and preparation, can compromise the performance of predictive models. Understanding the intricacies of data extraction and preparation challenges is essential for developing robust predictive models that can be reliably integrated into clinical practice.
Structured data quality assessments (Lewis et al., 2023; Miao et al., 2023; Thuraisingam et al., 2021; Nobles et al., 2015) are facilitated by frameworks like Weiskopf and Weng (Weiskopf and Weng, 2013) and METRIC (Schwabe et al., 2024). However, researchers often conduct unstructured assessments, documenting “lessons learned” from their projects (Ferrao et al., 2016; Sauer et al., 2022; Arbet et al., 2021; Maletzky et al., 2022; Sendak et al., 2017; de Kok et al., 2024). These assessments tend to focus primarily on data cleaning rather than the entire data extraction and preparation process, including cohort definition, outcome definition, and feature engineering. While literature on data quality assessments exists, mitigation strategies remain largely undocumented (Honeyford et al., 2022).
We identified and categorized the challenges encountered during EHR data extraction and preparation for predictive modeling. We provided actionable recommendations that can enhance data quality and improve the real-world applicability of dynamic prediction models in clinical settings. By addressing these challenges, we hope to contribute to the development of more reliable and effective predictive modeling frameworks that can ultimately benefit patient care.
@@ 2 Objective
This article provides a comprehensive list of challenges encountered during data extraction and preparation using electronic health record (EHR) data for developing dynamic prediction models for clinical use. It further proposes recommendations with actionable insights, intended for improving the quality of research and the practical applicability of clinical predictive models. Our insights are drawn from a selective literature review, as well as our experience with various EHR data extractions. This list is intended as a hands-on resource for data extraction engineers (who perform the data extraction) and researchers (who prepare the data for model building) to consult during the extraction and preparation process.
We focus on single-hospital structured data, covering both ICU-specific and hospital-wide extractions. We focus on medium to large-scale extractions, i.e. generated through structured ETL processes for a large number of extracted items spanning diverse clinical domains (laboratory results, medications, demographics, comorbidities, etc.), and we cover standardized (e.g., OMOP CMD) and non-standard, private extractions. We do not address combined data registries, such as national registries (integrating surveys, general practice, insurance data, and EHR extractions), multi-center data extractions, hospital data collected for clinical trials, or the extraction and processing of unstructured data (e.g., text notes or images). We cover the broader application of dynamic models (or continuous prediction) for offering the most comprehensive framework, although many of the recommendations are applicable to static prediction models (e.g., a single prediction per admission at 24 hours after admission).
@@
The section “EHR Data Flow from Data Collection to Model Building” explains the typical trajectory of data, for both model building and clinical implementation, setting the stage and providing context for the detailed list of challenges that follows. The section “Challenges and Recommendations” lists the challenges and actionable recommendations, which represent the core of the article. These are categorized into four groups: cohort definition, outcome definition, feature engineering, and data cleaning. This is one possible categorization, inspired by stages of data preparation process. The “Discussion and Conclusion” section summarizes the recommendations and reflects on their broader implications.
@@ 3 EHR data flow from data collection to model building
The data flow from the patient’s bed to the prediction model building typically follows three stages: data collection (a), data extraction (b) and data preparation (c) (Figure 1, I.). Patient data are either manually entered in EHR software modules or collected by devices and stored in one or multiple databases. The data collected serve multiple purposes, e.g., daily bedside clinical work, national benchmarking (van de Klundert et al., 2015), or reimbursing the care delivered (a). EHR data are then extracted from relational databases in a simplified (denormalized) format and typically stored in a data warehouse as multiple “base tables” per category, each capturing different aspects of patient data and healthcare events. Examples of base tables naming as per OMOP CDM are: PERSON, DRUG_EXPOSURE, DEVICE_EXPOSURE, CONDITION_OCCURRENCE, MEASUREMENT, NOTE, OBSERVATION, but naming and granularity of extraction can vary for non-standard extractions (b). Researchers building a prediction model either have access to the data warehouse or get a copy of all tables or, under specific ethical and legal considerations, a subset of the base tables or a subset of the patients in the data warehouse (e.g., admissions during a specified period of interest for patients undergoing mechanical ventilation). Further, they process the extracted data and bring it in a format on which a prediction model can be built (c).
@@
At model implementation time in clinical practice in the EHR software (Figure 1, II.), a trigger (e.g., update of lab results) or a scheduled task (e.g., every 24 hours) will initiate the request for a prediction. Data are already collected (a) for the patient in the EHR database(s). The same logic as for extraction (b) is reproduced (typically by reusing the queries or code used for extraction). Using the data packed in a specific format (e.g., json, xml), the prediction service (typically using a REST API interface for communication) is invoked. Data exchange between the EHR and the prediction service is generally performed using the FHIR (Fast Healthcare Interoperability Resources) standard. The prediction service will further prepare the data (c), invoke the model and return a prediction (and additional information, if foreseen) to the EHR software which will present it to users in the form of alerts or patient flags within the patient’s chart. The process is typically logged in the EHR system for monitoring purposes. In summary, at real-time prediction, the data collection (a) happens implicitly and is part of the normal clinical flow; data extraction (b) and data preparation (c) are identically reproduced as for model building.
@@
Figure 1: Data flow for model building pipeline (I.) and model implementation (II.) Two databases are exemplified as data sources, EHR and ICU, although multiple other sources might be used in the hospital’s flow and for data extraction. EHR = Electronic Health Record; DB = database; ICU = Intensive Care Unit; REST API = RESTful application programming interface
@@
We provide additional background on some particularities for each of the processes of collection, extraction and preparation, which represent transition phases from one data format to another:
@@
    Data collection: The EHR database will not reflect with maximum accuracy the “true state” of the patient. First, it will suffer of incompleteness, as not all possible markers and observations can be collected for all patients at all times. The decisions with regards to what data are collected (e.g., which laboratory tests are ordered and performed) is highly dependent on the patient conditions and on the hospital procedures. Data collected during routine clinical practice are generally documented more carefully when also used for national reporting of quality of care indicators. (van de Klundert et al., 2015) Sufficient data are though collected to support the patient’s clinical follow-up and treatment. From this perspective, EHR data differ vastly from data collected for clinical trials, where researchers specify the measurements, measurement methods and collection procedure. Second, nurses and clinicians might have slightly more information than the data collected in the system, either from patient conversations or organizational knowledge. Considering that we do not cover extraction of text notes and reports, tabular data only will always suffer of a level of incompleteness. Nevertheless, tabular data might prove sufficient for specific prediction tasks. Third, both manual data entering and data collected by devices can be at times error prone, software can have bugs and data recording procedures in the system will affect the granularity of observed data and will change over time. These considerations, which can be summarized as data completeness, correctness and currency, as defined by Weiskopf and Weng (Weiskopf and Weng, 2013), have to be carefully considered by researchers to assess if EHR data are fit for the prediction goal (Thuraisingam et al., 2021).
@@
    Data extraction: The extracted data might not correctly reflect the EHR database. Although data extraction generally aims for completeness of all clinically relevant data, the extraction process can introduce undesired artifacts or can have its own limitations. Whenever extraction logic bugs are detected, these should be corrected and safeguarded through unit and/or integration tests. Mature extraction platforms, tested through repeated use and proven reliability, will generally be less error prone and researchers can consider the maturity of the extraction platform as a factor influencing the time they will spend on data preparation. The extraction process is typically carried out by a data extraction engineer, data integration developer, data warehouse engineer or ETL developer representing the EHR vendor or the hospital IT department. Hereafter, we refer to this role as the data extraction engineer. They work closely with EHR software developers, hospital IT staff, and clinical personnel to understand database structures and data recording procedures. They perform clinical concept and terminology mapping and document the extracted data.
@@
    Data preparation: The prepared data might not correctly reflect the extracted data. Undesired artifacts can be introduced by feature engineering or data cleaning. Good documentation of the extraction format and close collaboration between researchers, data engineers and clinical experts will ensure that the data and the features are not misinterpreted. Coding errors can always occur. Time-sensitive data poses an additional challenge to ensure no temporal leaks (using future data to predict past events) are introduced by mistake. Outcome leaks (including outcome information in predictors) and test-train leaks (including information from test patients in the training datasets) can also occur if data are not preprocessed carefully. A good practice is to separate train and test sets first, and apply the data preprocessing separately. A modular organization of the code will facilitate unit testing (testing individual functions of a program in isolation to ensure they work as expected) and easy modifications without introducing new errors. The more complex the data preparation, the more error prone it becomes. The data preparation is performed by a researcher, data scientist, statistician or machine learning engineer, which we will refer to as researcher for the remainder of the paper.
Each stage of processing the data has its own challenges and can introduce new problems, widening the gap between the patient’s state and the data used by the prediction model and ultimately impacting the model building (suboptimal model), model evaluation (misleading performance metrics) or model implementation in clinical practice.
@@ 4 Challenges and Recommendations
We list common problems originating in the data collection process (a), the data extraction process (b) and the data preparation process (c). We provide recommendations for mitigation strategies that can be implemented during the data extraction (b) or data preparation (c). We also focus on problems that can impede the identical reproduction of the extraction and preparation at clinical implementation time. We have categorized the challenges and recommendations into four groups: (1) cohort definition (and inclusion/exclusion criteria), (2) outcome definition, (3) feature engineering, and (4) data cleaning, and each group contains problems originating in the collection, extraction or preparation process.
We provide mapping of the items to both the Weiskopf and Weng framework (Weiskopf and Weng, 2013) and the METRIC framework (Schwabe et al., 2024), whenever applicable. Weiskopf and Weng do not include dimensions covering the cohort representativeness and completeness of the extracted features, which are important in prediction settings. We will use the Completeness dimension to refer to completeness of data values, as in the original definition (Weiskopf and Weng, 2013), as well as completeness of the cohort and of the extracted features (our extension).
While we aimed to make these challenges as generic as possible, we recognize that specific issues are often unique to individual projects and may not apply to all prediction tasks. We advise users to assess the impact of each listed challenge in their specific context. Similarly, the recommendations might not always be universally applicable and depend on the project context. Our guidance remains pragmatic; at times, the best approach may be to “leave-as-is” to avoid the risk of overcorrection, which can backfire. While certain corrections may improve alignment with clinical meaning for a specific patient group, they can introduce errors or biases when applied to all patients or be difficult to reproduce at clinical implementation time, affecting the model’s performance in clinical use. Following the fitness-for-use principle (Miao et al., 2023), we encourage readers to remain pragmatic and address only the issues relevant to their data and prediction task.